### v1

# Path-integral Monte Carlo estimator for the dipole polarizability of quantum plasma

Abstract
We present a path-integral Monte Carlo estimator for calculating the dipole polarizability of interacting Coulomb plasma in the long-wavelength limit, i.e., the optical region. Unlike the conventional dynamic structure factor in reciprocal space, our approach is based on the real-space dipole autocorrelation function and is suited for long wavelengths and small cell sizes, including finite clusters. The simulation of thermal equilibrium in imaginary time has exact Coulomb interactions and Boltzmann quantum statistics. For reference, we demonstrate analytic continuation of the Drude model into the imaginary time and Matsubara series, showing perfect agreement with our data within ranges of finite temperatures and densities. Method parameters, such as the finite time-step and finite-size effects prove only modestly significant. Our method, here carefully validated against an exactly solvable reference, remains amenable to more interesting domains in higher-order optical response, quantum confinements and quantum statistical effects, and applications in plasmonics, heterogeneous plasmas and nonlinear optics, such as epsilon-near-zero materials.
preprint: APS/123-QED
I Introduction
The Drude model, introduced at the beginning of the last century by Paul Drude as an attempt at a comprehensive electron theory of metals [1, 2], remains nowadays ubiquitous in describing the optical response of metals, as well as optical materials with metal-like behaviours, such as epsilon-near-zero (ENZ) media [3]. The Drude parameters have been fitted to experiments [4] and hydrodynamic models of various kinds typically recover the Drude model or something close to it [5]. One of the advantages of the Drude model, beside its simplicity, resides in its high effectiveness in describing the optical response in a broad range of frequencies, from the near IR to the UV regime. This allows to characterize the behaviour of different photonic materials, such as metals [6, 7], plasmonic circuits [8], ENZ materials [9], as well as more complicated photonics platforms, such as metasurfaces [10], with an effective Drude model.
Thus, there is great intrigue towards the recovery of the effective response properties from first principles. In spirit, the Drude model describes the dynamics of non-interacting free electrons, which occasionally scatter from ions. When the scattering is rare, this model becomes the homogeneous electron gas (HEG), or jellium, whose quantum mechanical linear response function is the Lindhard function $\chi_{0}(\mathbf{q},\omega))$ [11], where $\mathbf{q}$ and $\omega$ treat, respectively, the momentum and the frequency of the incoming field. Not surprisingly, the Lindhard function, despite begin famously hard to evaluate generally, recovers the Drude model upon taking approximations of long wavelength, high temperature and low scattering [7, 6].
The Lindhard response can be expressed more generally as [12, 13]
$\chi(\mathbf{q},\omega)=\frac{\chi_{0}(\mathbf{q},\omega)}{1+v_{q}(1-G(\mathbf{q},\omega))\chi_{0}(\mathbf{q},\omega)},$ (1)
where the fraction treats dielectric screening, $v_{q}=e^{2}/q^{2}\varepsilon_{0}$ is the reciprocal Coulomb potential (in atomic units) and $G(\mathbf{q},w)$ is the so-called local field correction, basically treating effects due to Fermion exchange and dynamic correlations. For decades, numerous works (e.g. Refs. [12, 14, 13]) have discussed and featured either static $G(\mathbf{q},\omega=0)$ or dynamic $G(\mathbf{q},\omega)$ in pursuit of the non-trivial response within the so-called regime of warm, dense matter (WDM).
Recently, significant progress has been made to calculate the dielectric response of the WDM with the path-integral Monte Carlo method (PIMC) [15, 16, 17, 18, 19, 13, 20, 21]. The PIMC approach can treat accurate quantum many-body interactions at non-trivial temperatures, and it is in principle only subject to numerical challenges due to the Fermion sign problem (FSP) [22, 23]. Within this framework, moreover, the dielectric response is directly connected to the dynamic structure factor through the intermediate scattering function [11]. The latter can be straightforwardly estimated from the PIMC walker as the density-density correlation function in the imaginary time. This function can be then analytically continued in the complex plane to recover the full dynamic response of the system. The inverse transformation back to the real time, however, faces serious challenges with numerical noise, prevalent in Monte Carlo simulation, and thus requires special methods [24, 25, 20].
Within this framework, Dornheim et al. have comprehensively studied the first [13] and higher order [17, 18] response properties of the HEG. An apparent limitation of their approach is that the allowed values for the wavenumber are limited by $q=2\pi n/L$, where $n\geq 1$ and $L$ is the (cubic) simulation cell size. Therefore, the typical optical wavelengths with $q\ll L^{-1}$, thus requiring large $L$, are hard to accomplish because of the algorithmic scaling and performance loss attributed to the high particle count $N$. Indeed, the literature focuses on characterizing inelastic neutron scattering or X-ray Thompson scattering spectra, where the momentum transfers are significant [13] compared to that of photons.
Here, we explore a complementary PIMC approach to regain focus on the long, optical wavelengths. Namely, we demonstrate computation of the conventional optical response based on the complex dipole polarizability, or the dipole-dipole correlation function. This approach, which evades the difficulty of low $q$, has been studied before in finite systems [26, 27, 28, 29] but not in periodic ones. Granted, formal problems arise that are more naturally addressed in the Fourier space: The long-wavelength limit is only quasi-accurate, since we treat optical fields that are locally constant inside a finite box, i.e. $\lambda\gg L$, and formally expand that by summing over infinite periodic images. In practice, there is no field in the simulation, and thus, the approximation remains justified in all parameter ranges in the intended scope: for example, the box size $L$ required for $N=16$ particles to simulate the metallic density of gold is over $300$ smaller than an ultraviolet wavelength $\lambda=200$ nm.
However, it is not obvious, a priori, how a PIMC estimator for the dipole polarizability must be laid out for periodic systems or how it responds to finite-size effects or converge parameters, such as the finite time-step. Therefore, it is the aim of this work to establish these matters in a controllable setting: We first discuss analytical correspondence between the Drude model and the PIMC polarizability, evaluated in the imaginary time. For convenience and consistency with the Drude model, we use distinguishable particles, i.e. the Boltzmann statistics, to evade the Fermion sign problem—for now. We then survey various methodological and physical parameters to assert that the correspondence holds, when the Drude scattering rate is low. Indeed, we show that biases are statistically non-distinguishable within $2\%$ accuracy; the PIMC polarizability matches the Drude model. Formally, this is well anticipated, but practically the lack of biases is a signal of a robust simulation method. The assessments of performance help us avoid pitfalls, establish good practices and assert validity for continuation studies.
This work is organised as follows: In Sec. II, we lay out the theoretical context of the estimator and the PIMC simulation. In Sec. III we overview the computational details. In Sec. IV we discuss the numerical results and in Sec. V we summarize the work and future prospects of this methodology.
II Theory
In this section, we first overview the linear optical response of quantum particles in the long-wavelength limit. We then consider analytic continuation of the Drude model into imaginary domain, where it meets the dipole moment autocorrelation function accessible to the PIMC method. Throughout this work, we use the Hartree atomic units: $\hbar=1/(4\pi\epsilon_{0})=m_{e}=1$.
II.1 Electric field response
The first-order local polarization $\mathbf{P}(\mathbf{r},\omega)$ of charged plasma responding to an incident electric field $\mathbf{F}(\mathbf{r^{\prime}},\omega)$ can be written as
$\mathbf{P}(\mathbf{r},\omega)=\varepsilon_{0}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\mathbf{F}(\mathbf{r^{\prime}},\omega),$ (2)
where $\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)$ is a complex linear response function, the electric susceptibility, in the local response approximation. Let us consider the long-wavelength limit, in which the field is oscillating with a wavelength $\lambda=2\pi c/\omega\gg L$, where $c$ is the speed of light and $L$ is a measure of longitudinal distance containing the interaction. Then, within $L$ we may consider a locally constant field
$\mathbf{F}(\mathbf{r^{\prime}},\omega)\rightarrow\delta(\mathbf{r^{\prime}}-\mathbf{r})\mathbf{F}(\omega),$ (3)
where $\delta$ is the Dirac delta function, inducing total mean polarization $P(\omega)\equiv\varepsilon_{0}\chi(\omega)\mathbf{F}(\omega)$ according to total mean susceptibility
$\chi(\omega)=\frac{1}{L^{d}}\int_{0}^{L}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\delta(\mathbf{r}-\mathbf{r^{\prime}})d\mathbf{r^{\prime}},$ (4)
where $d$ is the dimensionality of $\mathbf{r}$. We may further consider the per-particle dipole polarizability $\alpha(\omega)$ through
$\varepsilon_{0}\chi(\omega)=n\alpha(\omega),$ (5)
where $n=N/L^{3}$ is the number density of the particles within $L$. The tensorial characters of $\chi$ and $\alpha$ are omitted here for simplicity of notation; because of the radial symmetry of the plasma, we only consider the diagonal, where $\mathbf{P}$ and $\mathbf{F}$ are parallel.
For quantized plasma represented by $N$ particles, the per-particle dipole polarizability $\alpha(\omega)$ can also be expressed using the Kubo formula [30]
$\alpha(\omega)$ $=-\frac{i}{N\hbar}\int_{0}^{\infty}\mathrm{d}te^{i\omega t}\int d\mathbf{r}\langle[\mu(\mathbf{r},t),\mu(\mathbf{r},0)]\rangle$ (6)
$\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}G^{R}_{\mu\mu}(t)$ (7)
$\equiv-G^{R}_{\mu\mu}(\omega),$ (8)
where $G^{R}$ denotes the retarded dipole-dipole correlation function and the dipole moment operator is given by
$\mu(\mathbf{r},t)=\sum_{i=1}^{N}q_{i}[\delta(\mathbf{r}-\mathbf{r_{i}}(t))-\langle\mathbf{r_{i}}\rangle],$ (9)
where $q_{i}$ is the charge and $\mathbf{r_{i}}(t)$ the time-dependent displacement of the $i$th particle, and $\langle\mathbf{r_{i}}\rangle$ is their mean displacement, a non-perturbed reference point. Clearly, when $\sum q_{i}\neq 0$ the value depends on the reference point, which is problematic in translationally invariant systems, such as plasmas. Therefore, it is more meaningful to consider the relative correlation function
$\tilde{G}^{R}_{\mu\mu}(t)\equiv G^{R}_{\mu\mu}(t)-G^{R}_{\mu\mu}(0)$ (10)
which is zero at $t=0$ but nevertheless yields
$\tilde{\alpha}(\omega)$ $\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}\tilde{G}^{R}_{\mu\mu}(t)=-G^{R}_{\mu\mu}(\omega)=\alpha(\omega)$ (11)
for all $\omega\neq 0$, while $\alpha(\omega=0)$ remains elusive. On the other hand, it is well known that the total response of charged gas in the limits of $\omega=0$ and $q=0$ (long wavelength) is zero due to perfect screening [31]. The linear response framework considered here does not treat screening, but this should justify why $\alpha(\omega=0)$ is unimportant for the free gas.
The relative polarizability $\tilde{\alpha}(\omega)$ is analytic in the upper complex plane and can be analytically continued to imaginary time $it\rightarrow\tau$. The associated, time-ordered imaginary-time correlation function is given by
$G_{\mu\mu}(\tau)=\langle\mathcal{T}\mu(0)\mu(\tau)\rangle,$ (12)
where $\mathcal{T}$ is the time-ordering operator. The analytic continuation can be done via the spectral density function
$A(\omega)\equiv-2\mathrm{Im}[G^{R}_{\mu\mu}(\omega)]$ (13)
where the bosonic integral transformation yields
$G_{\mu\mu}(\tau)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega),$ (14)
where $G_{\mu\mu}(\tau)=G_{\mu\mu}(\tau+n\beta)$ for any integer $n$ and $\beta=1/k_{B}T$ is the inverse temperature. It is easy to see that
$\tilde{G}_{\mu\mu}(\tau)$ $\equiv G_{\mu\mu}^{R}(\tau)-G_{\mu\mu}^{R}(0)$ (15)
$=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1-\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega).$ (16)
Alternatively, we may write
$G_{\mu\mu}(i\omega_{n})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{i\omega_{n}-\omega}A(\omega),$ (17)
where $\omega_{n}=2\pi n/\hbar\beta$ are the bosonic Matsubara frequencies.
While the imaginary-domain functions $G_{\mu\mu}(\tau)$ and $G_{\mu\mu}(i\omega_{n})$ contain the same information as $G^{R}_{\mu\mu}(\omega)$, the conversion from imaginary to real domain by inverting Eqs. (14) or (17) is a known ill-posed problem, which is accentuated by the unavoidable error bars present in Quantum Monte Carlo methods [24]. The use of existing numerical methods of analytic continuation, such as Maximum Entropy (MaxEnt) [25, 26], will be left for future consideration.
II.2 Analytic continuation of the Drude model
The ubiquitous model for considering the optical response of microscopic particles is the Drude model. The Drude susceptibility is given by
$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (18)
where $\omega_{p}^{2}=\tfrac{ne^{2}}{m\varepsilon_{0}}=\tfrac{e^{2}}{4\pi\varepsilon_{0}m}\tfrac{3}{r_{s}^{3}}$ is the plasma frequency and $\Gamma$ is a positive damping coefficient. The Drude per-particle polarizability is
$\alpha^{D}(\omega)=\frac{\chi_{e}^{D}(\omega)\varepsilon_{0}}{n},$ (19)
whose spectral density, according to Eq (13), is given by
$A^{D}(\omega)$ $=\frac{2\omega_{p}^{2}\varepsilon_{0}}{n}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}$ (20)
$=\frac{2e^{2}}{m}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}.$ (21)
The spectrum diverges at $\omega=0$ and its analytic continuation with Eq. (14) becomes problematic. However, the relative polarizability $\tilde{\alpha}(\omega)$ can be obtained from Eq. (16) because the divergence is controlled at $\omega=0$ by the transformation kernel.
Furthermore, an analytical expression for $G^{D}_{\mu\mu}(i\omega_{n})$, which is the correlation function in Matsubara frequencies associated to the Drude model, can be obtained by using Eq.  (20) and substituting into Eq. 17. This yields
$G^{D}_{\mu\mu}(i\omega_{n})=\frac{2\pi e^{2}}{m}\left[\frac{1}{\omega_{n}^{2}/\Gamma+\omega_{n}}\right].$ (22)
The analytical expression given by Eq.  (22) is valid for $\omega_{n}>0$. The static limit is defined here based on the zero-frequency Fourier transform of relative correlation function
$\tilde{G}^{D}_{\mu\mu}(i\omega_{n})=\int_{0}^{\beta}d\tau\tilde{G}^{D}_{\mu\mu}(\tau),$ (23)
but as discussed before, it is not physical and only used for numerical assessment of performance.
From Eq. (20) it is clear that the Drude spectrum does not depend on the physical density, because the oscillators are not interacting; rather, the density only shows in the transformation to the macroscopic susceptibility with Eq.(5). To consider the interacting problem, let us therefore turn to numerical path-integral simulation.
II.3 PIMC estimator
The PIMC method can be used to simulate interacting quantum particles at finite temperatures. The method draws samples from the distribution of the thermal density matrix:
$\rho(R,R;\beta)=Z^{-1}\int dR\langle R|\hat{\rho}(\beta)|R\rangle,$ (24)
where $R$ denotes a set of variable particle coordinates, $\hat{\rho}(\beta)=e^{-\beta H}$ is the density operator at inverse temperature $\beta$, $H$ is the Hamiltonian and $Z=\langle R|\hat{\rho}|R\rangle$ is the thermal partition function. Equation (24) overlooks summations arising from spin degrees of freedom and particle permutations required to treat proper quantum statistics of bosons or fermions; instead, the simulation treats so-called boltzmannons, which comprise a Coulomb interacting gas of distinguishable quantum particles. The boltzmannons evade the FSP but do not correspond to physical electrons, alas, except at high temperatures and low densities [23] (or strong coupling [32]), where particle repulsion outweighs exchange and the quantum statistics merge. Our results are also affected by the (lack of proper) quantum statistics.
The interacting Hamiltonian for $N$ distinguishable particles is given by
$H(R)=T+V=\lambda_{m}\sum_{i}\nabla_{i}^{2}+\sum_{i<j}\frac{q_{i}q_{j}}{|\mathbf{r}_{i}-\mathbf{r}_{j}|}$ (25)
where $\lambda_{m}=1/2m_{e}$ and $T$ and $V$ correspond, respectively, to the summation over kinetic and potential operators, the last ones comprising exact Coulomb interaction between pairs of particles. Since $T$ and $V$ do not commute, the density operator must be expanded with the semi-group property
$e^{-\beta(T+V)}=\lim_{M\rightarrow\infty}\left[e^{-\Delta\tau T}e^{-\Delta\tau V}\right],$ (26)
where $\Delta\tau=\beta/M$ is the time-step and $M$ is the Trotter number. Using Eq. (26) and the product property [23], the exact Coulomb action between pairs of species can be treated for finite time-steps [33]. In this so-called pair approximation, higher-body interactions retain a modest time-step error proportional to $(\Delta\tau)^{3}$ [34, 33]. Therefore, it is justified to choose finite $M$, making the PIMC integrand a discrete walker of $M$ times $R$ coordinates, which forms a closed loop, an imaginary-time trajectory:
$\rho(R,R;\beta)=Z^{-1}\prod_{j=1}^{M}dR_{j}\langle R_{j-1}|\hat{\rho}(\Delta\tau)|R_{j}\rangle,$ (27)
where $R_{j}$ contains the real-space coordinates for $N$ particles at $j$th time-slice, and the loop closes with $R_{M}=R_{0}$. The trajectory is efficiently sampled as a Markov chain using the Metropolis multilevel bisection algorithm, as described in [35, 23, 33].
The calculation of the thermal expectation values of diagonal properties, such as potential energy $V$, is done from Eq. (27) as
$\langle V\rangle$ $=Z^{-1}\int dR\langle R|\hat{V}\hat{\rho}(\beta)|R\rangle$ (28)
$=Z^{-1}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)V(R_{0})$ (29)
$=\frac{1}{M}\sum_{j=1}^{M}\langle V(R_{j}),$ (30)
where the last line employs symmetry properties of the trace, enabling to calculate mean over all time-slices of any trajectory. Because of the kinetic operator, the total energy $E$ is more complicated; here we use the thermal estimator [23], whose numerically exact pair energy $\partial\rho/\partial\beta$ can be precomputed [33].
The calculation of time-ordered observables, such as the imaginary-time correlation function from Eq. (12), is equally straightforward. The mean per-particle correlation function is evaluated as
$\langle G_{\mu\mu}(\tau)\rangle=\frac{1}{ZN}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)\mu(R_{0})\mu(R_{t}),$ (31)
where the time-difference index is $t=j+\tau/\Delta\tau$, modulo $M$. The expression can be easily symmetrized, optimized for evaluation and generalized for different combinations of measurement operators, as detailed in [28, 29]. In the path-integral context, the dipole moment operator is
$\mu(R_{t})$ $=\sum_{i=1}^{N}q_{i}\left[\mu(\mathbf{r}_{i}(t\Delta\tau)-\langle\mu_{i}\rangle\right],$ (32)
which again depends on the point of reference $\langle\mu_{i}\rangle$. However, as was done in Ref. [26], we identify
$\langle\tilde{G}_{\mu\mu}(\tau)\rangle$ $=\frac{1}{2N}\langle(\mu(0)-\mu(\tau))^{2}\rangle,$ (33)
where the relative correlation function arises from mean squared fluctuations of the instantaneous dipole moment along the imaginary-time path.
II.4 Periodic boundary conditions
The simulation of charged plasma requires periodic boundary conditions (PBC). Here we use a cubic box of size $L$, whose size is determined by
$L=r_{s}\left(\frac{4\pi N}{3}\right)^{(1/3)},$ (34)
where $r_{s}=(3/4\pi n)^{1/3}$ is the Wigner-Seitz radius of the particle gas. The treatment of periodic Coulomb images is treated using the ordinary Ewald summation [36, 37]. Note that the long-wavelength approximation discussed in Sec. II.1 also uses $L$, and the symbols are in practice interchangeable but semantically different.
The finite density is enforced for the many-body interactions of the plasma, but otherwise the evaluation resembles an open-boundary problem. So, the evaluation of the dipole polarizability estimator, Eq. (33), is done on walkers whose all particle paths are unfolded into continuous trajectories near boundaries. Since the simulated particles may not exchange paths (permute), non-zero winding of a path through the cell is an artifact of the finite size. Therefore, we reject all moves leading to non-zero winding. Beside this constraint, the sampling is done using the multilevel bisection method [23]. The probability of winding increases when the thermal wavelength of the bisection move approaches the cell dimension, i.e.
$\sqrt{2^{l}\lambda_{m}\Delta\tau}\sim L,$ (35)
where $l$ is the bisection multilevel. We constrain $l$ so that the l.h.s. of Eq. (35) is lower than $L/8$. This is a simplification to avoid using periodic free-particle density matrix [23, 33], which would not inhibit the winding, anyway. As a result, fewer than one winding move per $10^{3}$ trial moves is observed and has to be rejected in any density. A modest sampling bias is mitigated along other finite-size effects when $L$ naturally grows with $N^{1/3}$.
III Computational details
The PIMC simulation software, including the exact Coulomb pair action generation, implements the canonical Metropolis Monte Carlo algorithm for a PIMC walker in Fortran90, featuring parallel sampling and data binning of a number of Markovian walkers. Here, 40 CPU cores are used for parallel sampling. Raw data after sample binning is stored in hierarchical data format (HDF5) and data processing chains operate on standard numerical Python libraries. The integral transformations, such as Eq. (16), are carried out using numerical quadrature. The simulation campaigns are operated with Nexus workflows [38] on various high-performance computing (HPC) facilities listed in Sec. V. All implementation details, including code snapshots and workflow scripts for step-by-step reproduction of the simulations, postprocessing and individual figures are supplied in a separate data repository [39].
[…]
IV Results
Figure 1: Comparison overview of the relative dipole correlation function $\tilde{G}_{\mu\mu}$ from PIMC and the Drude model. In the top, the imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ is plotted at two temperatures, $\Theta_{1}=0.1$ and $\Theta_{2}=0.2$, resulting in two curves ending at different values for $\beta_{1}=2\beta_{2}$. In the bottom, the first few corresponding Matsubara data $\tilde{G}_{\mu\mu}(i\omega_{n})$ are plotted, where the frequencies $\omega_{n}=2\pi n/\hbar\beta$ for $\Theta_{1}$ are twice those of $\Theta_{2}$, but the two functions lie on the same curve. In both cases, the data match the analytically continued Drude model within statistical uncertainties.
Let us now discuss results of the PIMC simulation, laid out in the earlier sections, as applied to charged quantum plasma at various finite temperatures and densities. With $m=m_{e}$ and $q^{2}=e^{2}$, the simulated particles resemble free electrons, only without sign-flipping permutations. Thus, we use here conventional units of the HEG and express densities with $r_{s}$ and finite temperatures with $\Theta=T/T_{f}$, where $k_{B}T_{f}=\hbar^{2}k_{f}^{2}/2$ and $k_{f}=(3n\pi^{2})^{1/3}$ are, respectively, the Fermi temperature and the Fermi wavevector. The covered densities range between $r_{s}=2.0\ldots 8$ a0 and finite temperatures between $\Theta=T/T_{f}=0.1\ldots 0.8$.
The main observables of interest are overviewed in Fig. 1: the relative imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ and its Matsubara series $G_{\mu\mu}(i\omega_{n})$. As stated in Sec. II, the vertical offset of $G_{\mu\mu}(\tau)$ is arbitrary, and thus, we treat it as zero, making $\tilde{G}_{\mu\mu}(\tau)<0$ for all $\tau$. Again, the static Matsubara frequency $\omega_{n}=0$ must necessarily come out negative:
$G_{\mu\mu}(i\omega_{n}=0)=\int_{0}^{\beta}d\tau\tilde{G}_{\mu\mu}(\tau)<0,$ (36)
which stands out from the the rest, $G_{\mu\mu}(i\omega_{n}>0)>0$. Therefore, for convenience we will treat it with the opposite sign in figures and analyses, namely $\tilde{G}_{\mu\mu}(i\omega_{n}=0)\equiv-G_{\mu\mu}(i\omega_{n}=0)$. According to Eq. (11), $G_{\mu\mu}$ is directly connected to the physical polarizability for finite frequencies, i.e., $\alpha(\omega\neq 0)$. However, we will not seek here the ill-posed analytic continuation from imaginary to real domain.
Rather, for validation of the results for $\tilde{G}_{\mu\mu}$, we compare the PIMC data to the Drude model, given in Eq. (22), in the imaginary domain. Thus, we report the relative error $\Delta\tilde{G}_{\mu\mu}=(G_{\mu\mu}-G^{D}_{\mu\mu})/G^{D}_{\mu\mu}$ for both the corresponding autocorrelation and Matsubara functions. Unless stated otherwise, we use $\Gamma=0.0001$ in the Drude model to approximate low-$\Gamma$ limit while maintaining numerical stability. As will be seen, the correspondence is close and $\Delta\tilde{G}_{\mu\mu}\neq 0$ is rarely statistically significant. Therefore, for the first-order proxy of accuracy we will assume
$\eta\equiv\Delta\tilde{G}_{\mu\mu}(i\omega_{n}=0),$ (37)
which captures principal scaling differences between the PIMC and the Drude model, if any should be found.
Figure 2: Scaling of per-particle kinetic energy $K/N$ (left) and potential energy $V/N$ (bottom) with finite NN and variable dimensional Ewald cutoff parameter $r_{c}k_{c}$. The solid lines correspond to raw PIMC observables and the dotted lines are finite-size corrected values based on Eqs. (38) and (39). The case $r_{c}k_{c}=0$ means no analytical Ewald summation of any kind, including neutralizing background, and so the $E_{p}$ is off above the plot.
Let us considering systematic merits of accuracy for the PIMC simulation. Finite-size effects of the energies are studied by varying the number of particles $N$ in the simulation, while keeping the density fixed. Results for kinetic and potential energies are presented in Fig. 2 for up to $N=16$ and $\Theta=0.2$ at $r_{s}=4$. Shown are also finite-size corrections [40, 34] ($\xi=0$), which rectify the energies toward the thermodynamic limit (TDL), according to:
$K_{p}$ $=\frac{K_{N}}{N}+\frac{1}{N}\left(\frac{\omega_{p}}{4}+\frac{5.264}{4\pi r_{s}^{2}(2N)^{1/3}}\right)$ (38)
$V_{p}$ $=\frac{V_{N}}{N}+\frac{\omega_{p}}{4N},$ (39)
where the corrected energies $K_{p}$ and $V_{p}$ are further multiplied with $\mathrm{tanh(\omega_{p}\beta)}$. The correlation energy is getting saturated with $N$, which also influences the kinetic energy. Simulation with $N>16$, or a reliable extrapolation to the TDL, would be needed to resolve the statistically significant finite-size bias of the energies. Also, clearly the Ewald summation with $k_{c}r_{c}>10$ is required to eliminate statistically significant bias from treating periodic images.
Figure 3: Scaling of $\tilde{G}_{\mu\mu}(\tau)$ (left) and $\tilde{G}_{\mu\mu}(i\omega_{n})$ (right) with finite NN. The top panels show absolute values of the PIMC data (solid) and the corresponding Drude value (dotted). The bottom panels shows the relative differences $\Delta\tilde{G}_{\mu\mu}(\tau)$ (solid lines) and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$, including statistical uncertainties (dotted lines).
To make physical benchmarks, there is obvious systematic bias related to the quantum statistics, i.e., the lack of Fermion exchange in the simulation. Otherwise, the Coulomb pair-density matrix is numerically exact and there are no nodal constraints. Remaining bias ensues from using a finite time-step $\Delta\tau$ to treat higher-body correlations (beyond pair approximation) and periodic images. As the higher-body correlations intensify with density, we assume for this work a choice of the time-step according to $\Delta\tau=ars^{3}$, where $a=0.001$ Ha-1.
The finite-size effects of the dipole polarizability are more topical: The scaling of $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $N$ is plotted in Fig. 3 for $k_{c}r_{c}=15.0$. Overall, the effects of $\tilde{G}_{\mu\mu}$ from PIMC versus Drude due to finite size are negligible regardless of $N$. Nevertheless, we will use $N=16$ for the remainder of this work. The effects due to finite time-step and the Ewald cutoff parameter $k_{c}r_{c}$ are also statistically insignificant.
The lack of finite-size effects in $\tilde{G}_{\mu\mu}$ is in stark contrast with the energies, which are affected by the correlation and, although absent here, exchange. The indifference to the number of particles is therefore a manifestation of a greater principle: the trajectories of an ensemble of free particles in a periodic setting sample the phase space in collective harmony. The dipole autocorrelation estimator only measures the collective state and distributes it evenly for each particle. Another way of seeing this is that the low-momentum states of the plasma are occupied and cannot be used to excite the dielectric response of individual particles. Curiously, the cases of interacting (PIMC) and non-interacting (Drude) come out effectively the same.
Figure 4: On the left, $\tilde{G}_{\mu\mu}(\tau)$ and $\Delta\tilde{G}_{\mu\mu}(\tau)$ with variable $\Theta$ at $r_{s}=4$. On the right, $\tilde{G}_{\mu\mu}(i\omega_{n})$ and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$ with variable $\Theta$ at $r_{s}=4$. Solid lines represent the absolute data and dotted lines the statistical uncertainties.
Finally, we study how $\tilde{G}_{\mu\mu}(\tau)$ depends on physical parameters, the finite temperature and density. In Fig. 4 we plot $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $\Theta$ with $r_{s}=4$ and $N=16$. The plotting axes of $\tilde{G}_{\mu\mu}(\tau)$ are rescaled to $\tau/\beta$, emphasizing that the absolute values of $\tilde{G}_{\mu\mu}(\tau=\beta/2)$ decrease with the inverse temperature. Consequently, the Matsubara integral $\tilde{G}(i\omega_{n}=0)$ goes down with $\beta$, but otherwise the Matsubara series $\tilde{G}(i\omega_{n}>0)$ follow the same curve, only with different spacings. Clearly, the PIMC results match the Drude model within statistical uncertainty of $2\%$ or less.
In Figure 5 we discuss how the match depends on the Drude relaxation parameter $\Gamma$. When $\Gamma$ is increased, the Drude correlation function $\tilde{G}_{\mu\mu}(\tau)$ loses magnitude. The phenomenological explanation is that the scattering rate of the charge carriers from ions increases with $\Gamma$, and thus, the spatial diffusion slows down. The PIMC simulation of free particles has no scattering, and thus, the correspondence is the best with low $\Gamma$. This is presented in the lower panel of Fig. 5, where linear fits approach $\eta\rightarrow 0$ when $\Gamma\rightarrow 0$. The relative error $\eta$ scales linearly with $\Gamma$ and $\beta$, and so, curves representing $\eta/\beta$ are overlaid within statistical uncertainties. With this consideration we want to suggest a future possibility of estimating $\Gamma$ from PIMC simulations, provided a reasonable but yet unknown sampling method of the scattering process. As seen from all the presented results, such a simulation could be easily validated, while providing first-principles insights to the Drude scattering. Furthermore, using a finite-$\Gamma$ simulation such as that could also lend itself for studying nonlinear optical phenomena of Drude-like materials on a quantum mechanical level.
Figure 5: On the left, a comparison of $\tilde{G}_{\mu\mu}(\tau)$ versus $\tau$ from PIMC (solid black line) and the Drude model with variable values for $\Gamma$. On the right, the rescaled goodness measure $\eta/\beta$ is plotted against $\Gamma$ based on variables temperatures and $r_{s}=4$. Dotted lines are linear fits to the data.
V Summary and outlook
We have investigated a PIMC estimator for the dipole polarizability of periodic quantum gases and plasmas in the long-wavelength limit, i.e., the optical regime. We have bench marked the estimator with quantum Coulomb plasma in a range of finite densities ($r_{s}=2\ldots 8$ a0) and finite temperatures ($\Theta=0.1\ldots 0.8$ and up to $T=10^{6}$ Kelvin). The PIMC results are compared to the Drude model in the imaginary domain, where they match within statistical limits and $\Delta\tilde{G}_{\mu\mu}<2\%$, when the Drude damping coefficient $\Gamma$ is small. That is, the PIMC results for $\tilde{G}_{\mu\mu}$ show no statistically significant effects due (reasonable) simulation parameters, such as finite time-step, periodic images or finite number of particles, nor physical parameters such as finite density and temperature. This corroborates that the Drude spectrum is suitable for describing the linear optical response (i.e. the long-wavelength limit) regardless of the coupling strength. Perhaps some minor methodological, systematic biases would have surfaced with generous statistical sampling, but this does not appear as a worthy goal.
In summary: due to the collective motion of many interacting particles in a flat potential, all results come out consistent, effectively, with a free particle. This physical conclusion may not be overly interesting, but it firmly establishes that the numerical method, including the complete post-processing and analysis framework, is valid and robust. To this end, we have demonstrated special sampling approaches (see Sec. II.4), to allow accurate simulation even with a low particle count. These practices form a baseline for numerous further developments:
i) The Fermion spectrum. The treatment of signed permutations should be considered in the future, to unfold the true electron gas behaviour, including complexities in the statistical performance, sampling acceleration and path winding.
ii) Confined spectrum. Treating interacting particles in environments less regular than a flat potential (e.g. weak confinements, optical cavities or ionic plasma) should introduce non-trivial thermal and quantum effects for example with respect to the Drude–Lorentz model.
iii) Inverse Laplace transform. Special methods to invert Eqs. (16) or (17) could be studied to migrate possible quantum corrections into effective models. The MaxEnt method is not the best suited for the broad, Drude-like spectra [23, 41] of this work, but in the presence of resonance peaks, it could be.
iv) Higher-order response properties. The response functions beyond the dipole order could be more sensitive to thermal quantum many-body effects. Their static and dynamic variants, such as $\chi^{(3)}$ of the Drude-like materials, could inform effective models [5] and applications, such as ENZ materials [42].
Acknowledgements.
The authors wish to acknowledge CSC – IT Center for Science, Finland, and the Tampere Center for Scientific Computing; for computational resources. The authors acknowledge the financial support from the Photonics Research and Innovation Flagship (PREIN - decision 320165) and the Research Council of Finland project AQUA-PHOT (decision Grant No. 349350). JT acknowledges Ilkka Kylänpää for his support in the implementation and assessment of periodic potentials.